OpenStudy (anonymous):

HELP!!!! The temperature on Wednesday was 8° more than a third of Tuesday’s temperature. If T represents Tuesday’s temperature, write an expression for Wednesday’s temperature, W.

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

@briannababe

OpenStudy (anonymous):

how you write 1/3 of T ?

OpenStudy (anonymous):

If you want to know 1/3 of something, do you add, subtract, multiply, or divide?

OpenStudy (anonymous):

I will put up my answer options. A. W = 3T + 8 B. W = 3T – 8 C. W=1/3T+8 D. W= 1/3T -8

OpenStudy (anonymous):

cool. so two of those of 3T and two have (1/3)T. What do you multiply T by if you want to know 1/3 of T?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

multiplying by 8 gives you 8 times T. Your problem says "a third of Tuesday’s temperature." Only 2 of your options have 1/3. So which can you eliminate?

OpenStudy (anonymous):

Well it sounds like A and B are irrelevant, so it's got to be C or D I'm thinking C.

OpenStudy (anonymous):

yes C

OpenStudy (anonymous):

Thanks so much.